Please help me, I don't understand how to do this at all-
"The volumes of two similar solids are 2197 m3 and 64 m3. The surface area of the larger one is 845 m2. What is the surface area of the smaller one?"


A. 64 m2
B. 320 m2
C. 80 m2
D. none of these
Thank you!!!!!!!!!!!!!




Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

845/2197:n/64
2197n=54080
n=24.62 m2..

Answer by Alan3354(69443)

"The volumes of two similar solids are 2197 m3 and 64 m3. The surface area of the larger one is 845 m2. What is the surface area of the smaller one?"
-------------------
Volume is a function of the cube of the dimenions.
2197 = 13^3 and 64 = 4^3, so it's 13 to 4.
Area is a function of the square of the dimensions.
--> 13^2 to 4^2 = 169 to 16
-----------
845/169 = A/16
Area = 80 sq meters
------------------
A. 64 m2
B. 320 m2
C. 80 m2
D. none of these